### v1

# Mixtures of Neural Network Experts with Application to Phytoplankton Flow Cytometry Data

Abstract
Flow cytometry is a valuable technique that measures the optical properties of particles at a single-cell resolution. When deployed in the ocean, flow cytometry allows oceanographers to study different types of photosynthetic microbes called phytoplankton. It is of great interest to study how phytoplankton properties change in response to environmental conditions. In our work, we develop a nonlinear mixture of experts model for simultaneous clustering and regression utilizing random-weight neural networks. Our model allows one to flexibly estimate how cell properties and relative abundances depend on environmental covariates, without the computational burden of backpropagation. We show that the proposed model provides superior predictive performance in simulated examples compared to a mixture of linear experts. Also, applying our model to real data, we show that our model has (1) comparable out-of-sample prediction performance, and (2) more realistic estimates of phytoplankton behavior.
Keywords: Flow Cytometry, Mixture of Experts, Model-based clustering, Phytoplankton, Random Weight Neural Networks
1 Introduction
Phytoplankton are photosynthetic microbes that reside at the bottom of the marine food chain. They conduct more photosynthesis than all flora on land (Field et al.,, 1998), acting as an invaluable component of Earth’s biological carbon pump (Ducklow et al.,, 2015; Boyd et al.,, 2019). Thus, studying phytoplankton behavior and distribution in the ocean is directly relevant for our understanding of carbon transport and, by extension, the earth’s biogeochemical cycle (Mouw et al.,, 2016; Wang and Moore,, 2012). Phytoplankton have faced a number of environmental pressures in recent years, including increased levels of carbon dioxide, subsequent ocean acidification, and higher average annual temperatures (Van de Waal and Litchman,, 2020). Some recent work predicts major changes in relative abundances (Flombaum and Martiny,, 2021) in response to environmental changes. However, it is still largely unknown how these trends will affect phytoplankton abundance and photosynthetic activity. Thus, understanding the response of phytoplankton to various environmental conditions constitutes an important line of work.
An increasingly popular method to measure phytoplankton attributes in the ocean is flow cytometry (Sosik et al.,, 2003). A fluid sample is passed through an instrument which measures the size and fluorescence for each particle. Then, the data analyst uses expert judgment – aided by software – to group these particles’ measurements into distinct types. This process is called gating and has traditionally been done manually, heavily relying on the analyst’s visual inspection of the data (Verschoor et al.,, 2015; BD Life Sciences,, 2023). In recent years, automatic gating algorithms have been in active development as an alternative to manual gating. However, most currently available gating algorithms are motivated by biomedical applications and are unsuitable for marine flow cytometry data. Our primary focus is to simultaneously cluster particles into distinct species, and to quantify each species’ response to changes in the ocean’s environmental conditions in a flexible manner, using a custom statistical mixture model.
We briefly review the literature on statistical models for flow cytometry that are most relevant to our work. Recent model-based statistical methods typically assume the data come from a mixture of real-valued statistical distributions. Pyne et al., (2014) and Lee et al., (2016) use finite mixtures of multivariate skew $t$-distributions with sample-specific random effects. Dundar et al., (2014) and Cron et al., (2013) use hierarchical Dirichlet process mixtures of Gaussian kernels to borrow information across samples. Soriano and Ma, (2019) develop a specialized Dirichlet process Gaussian mixture which accounts for inter-sample variation in both the component parameters and weights. Gorsky et al., (2024) develop a Bayesian nonparametric mixture of hierarchical skew normal kernels, a hierarchy which captures both sample-specific and inter-sample structure. Hejblum et al., (2019) develop a Dirichlet process mixture of skew $t$-distributions which accounts for temporal ordering of samples. However, none of these approaches use environmental covariates. The most relevant method, outlined in Hyun et al., (2023), applies a mixture of experts (Jordan and Jacobs,, 1994) to phytoplankton flow cytometry data, modeling both cluster means and probabilities as linear functions of environmental covariates. We build on this last method by developing a mixture of nonlinear experts.
The use of mixtures of experts to perform model-based clustering is supported by a rich literature which has developed over the past thirty years. Recently, Huang and Yao, (2012) showed that a semiparametric mixture of experts is identifiable under weak assumptions, and Nguyen et al., (2021) showed that a mixture of experts using a softmax gating function and location-scale kernels can provide arbitrarily accurate density estimates. Also, the use of a neural network in a mixture of experts has gained popularity in recent years. For example, Hadavandi et al., (2016) use a mixture of neural network experts in an agglomerative clustering scheme, and Liu et al., (2020) propose a mixture of neural network regressions in which all neural network experts share a common set of parameters. However, the latter approach does not use covariates to inform the cluster probabilities, and both are computationally intensive to train, requiring algorithms such as backprogagation or batch-size gradient descent. Our model also uses a mixture of neural network experts, but we randomly draw and then fix the hidden layer weights, thus bypassing the need for heavy computation. Huang et al., 2006a show that such random-weight neural networks can approximate any data-generating function with arbitrary accuracy and still generalize well to held-out samples. This technique has been applied to a variety of settings, including nonlinear regression (He et al.,, 2016), clustering (He et al.,, 2014; Ding et al.,, 2017), and mixture modeling (Husmeier and Taylor,, 1998). See Cao et al., (2018) for a more thorough review.
The method developed by Hyun et al., (2023) assumes that Gaussian cluster means and the logits of the cluster probabilities are linear functions of the covariates. However, we show that the nonlinear functions are more apt for modeling flow cytometry data collected over a large spatial region. For instance, microbe characteristics such as cell size have biological upper and lower limits which linear functions cannot predict effectively. Generally, phytoplankton species’ attributes and abundance likely do not respond to environmental changes monotonically. Also, certain pairs of covariates may have an interaction effect – for example, increased nutrient availability are known to improve or hinder a cell’s ability to perform biological functions. Since the functional forms for nonlinear relationships are difficult to prescribe prior to modeling, a flexible nonparametric approach such as ours is useful.
Here, we extend the approach of Hyun et al., (2023) in two ways. First, we use a random-weight neural network to expand the class of estimable regression functions. Such a model can more flexibly model cluster means and probabilities as functions of environmental conditions, thus allowing the user to fit a single model to flow cytometry data collected from a wide, heterogeneous range of ocean conditions. Second, we use the principal components of environmental covariates as regressors rather than the original covariates themselves. This vastly reduces the number of regressors and removes the strong intercorrelations in the original covariates, thus easing model interpretation.
The rest of this article is organized as follows: in Section 2 we describe the data. In Section 3 we describe the motivation behind the model and the structure of the model itself. In Section 4 we provide a simulation study which provides strong evidence towards the idea that the proposed model fits linear data as well as or better than a mixture of linear experts and fits strongly nonlinear data better than the linear model. In Section 5 we apply the proposed model to real flow cytometry data and uncover new insights into the responses of phytoplankton to changes in environmental conditions. Finally, in Section 6 we review our methods and results before outlining recommendations for future work.
2 SeaFlow cruise data description
We will treat marine flow cytometry data as a scatterplot-valued discrete time series $\{\boldsymbol{y}^{(t)}\}_{t=1}^{T}$, where each scatterplot $\boldsymbol{y}^{(t)}$ is an $n_{t}\times d$ matrix. We call $\boldsymbol{y}^{(t)}$ a cytogram. Each cytogram has rows $\boldsymbol{y}_{i}^{(t)}\in\mathbb{R}^{d}$ consisting of the $d$ measured optical properties of the $i$th particle observed during time interval $t$, for $i=1,\ldots,n_{t}$. The total number of particles observed is $N=\sum_{t=1}^{T}n_{t}$. Finally, cytogram $t$ shares covariate values $\boldsymbol{x}_{t}\in\mathbb{R}^{p}$, where $p\ll n_{t}$ for all $t$.
The flow cytometry data we analyze are from the Gradients 2 cruise collected by the Armbrust Lab (Ribalet et al.,, 2019) in the summer of 2017. This cruise traversed a large spatial region with heterogeneous environmental conditions, making it a good candidate to showcase the flexibility of our model to variation in the relationships between response data and covariates. Specifically, the cruise traveled between the North Pacific Subpolar and Subtropical gyres, collecting measurements of phytoplankton abundance and optical properties using the continuous-time flow cytometer known as SeaFlow (Swalwell et al.,, 2011). Each particle $\boldsymbol{y}_{i}^{(t)}$ from a given hour-long interval $t$ has $d=3$ measurements: (1) cell diameter, and (2) red fluorescence and (3) orange fluorescence, corresponding to the light intensities of emission spectra associated with chlorophyll and phycoerythrin pigments. Since all three dimensions’ original measurements are highly skewed, they are transformed using the natural logarithm (then linearly rescaled to be roughly between 0 and 8). In total, there are $083$-dimensional cytograms. However, we later created a twelve-hour lagged version of a covariate, which led us to remove the first 12 cytograms for a final count of 296. Additionally, some observations were censored at SeaFlow’s measurement limits; these points were simply removed prior to modeling. Lastly, an auxiliary measurement $\{c_{i}^{(t)}\}_{i,t}$, the inferred biomass of each particle, will be used. We use this cytogram and biomass data as a regression response.
Accompanying the response data, there are $33$ environmental covariates sourced from the Simons CMAP database (Ashkezari et al.,, 2021), accessed via the cmap4r R package (Hyun et al.,, 2025), and aggregated to match the hourly time resolution of the response data. These environmental covariate data include biochemical and physical attributes collected from satellites (and processed as long-term averages) and from on-board instruments, for a total of $p=37$ covariates over $T=296$ time points corresponding to unique hours. The vector of all $37$ covariate values at time $t$ is written as $\boldsymbol{x}_{t}$, and $\boldsymbol{X}\in\mathbb{R}^{296\times 37}$ is the covariate matrix with row $t$ equal to $\boldsymbol{x}_{t}$. For more details regarding data curation, see Hyun et al., (2023).
Figure 1: Principal components of the environmental covariates. The top panel shows the sample correlations between the first four principal components, latitude, and the covariates. The bottom-left panel shows the first principal component (PC1), as well as the most scientifically relevant covariates which are strongly correlated with PC1, plotted over time. We also overlay latitude as a dashed black line to show its similarity to PC1. The two vertical dotted lines at latitudes 31.9 N and 34 N represent the estimated boundaries of the North Pacific Transition Zone, a region where latitude-varying conditions change rapidly, hand-picked as the region where PC1 changes rapidly. In the bottom-right panel, we display a similar plot for PC3 and PC4. The full set of principal components is shown in Supplement A.
The 37 covariates are numerous and strongly intercorrelated, so we apply principal components analysis (PCA) to the covariate matrix. We first calculate the singular value decomposition of the covariate matrix $\boldsymbol{X}$ as
$\boldsymbol{X}=\boldsymbol{U}\boldsymbol{D}\boldsymbol{V}^{\prime}$
and then calculate the $q$-dimensional orthogonal projection of the covariate matrix $\boldsymbol{\Psi}\in\mathbb{R}^{T\times q}$ as
$\boldsymbol{\Psi}=\boldsymbol{X}\boldsymbol{V}_{\cdot,1:q},$ (1)
where $\boldsymbol{V}_{\cdot,1:q}$ are the first $q$ columns of $\boldsymbol{V}$. In all experiments, we choose $q$ to be the lowest value such that the proportion of variance explained by the first $q$ principal components is at least 0.95. For the dataset used in our analysis, we found $q=9$, which we use throughout this article.
It is worth examining the principal components $\{\boldsymbol{\Psi}_{\cdot,j}\}_{j=1}^{4}$ in some detail. For brevity we will adopt the shorthand of, for example, PC1 to denote the first principal component. As shown in Figure 1, the first principal component PC1 has a strong correlation with latitude. Since PC1 and latitude share strong associations with environmental conditions such as nutrient levels, temperature, and salinity, information about these conditions can be compressed into a single variable which summarizes latitudinal variation in the ocean. The second principal component PC2 is most strongly associated with wind speed, sea level anomaly, and iron levels, summarizing information relating to physical oceanographic characteristics. PC3 is related to physical ocean characteristics such as zonal geostrophic velocities and lagged sunlight (photosynthetically active radiation or PAR for short) time-lagged by 3 and 6 hours. Lastly, PC4 is comprised of PAR and its 9- and 12- hour lagged variants.
3 Methodology
We assume cytograms follow a mixture of multivariate Gaussian distributions
$f\left(\boldsymbol{y}_{i}^{(t)}\right)=\sum_{k=1}^{K}\pi_{k,t}\phi_{d}\left(\boldsymbol{y}_{i}^{(t)}\mid\boldsymbol{\mu}_{k,t},\boldsymbol{\Sigma}_{k}\right),$
where $\phi_{d}(\ \cdot\ |\boldsymbol{\mu},\boldsymbol{\Sigma})$ is the probability density function of a $d$-dimensional Gaussian random vector with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Introducing a latent membership $z_{i}^{(t)}\in\{1,\cdots,K\}$, we can write the distribution of particle $\boldsymbol{y}_{i}^{(t)}$ given latent membership $z_{i}^{(t)}$ as:
$\left(\boldsymbol{y}_{i}^{(t)}|z_{i}^{(t)}=k\right)$ $\sim\mathcal{N}_{d}\left(\boldsymbol{\mu}_{k,t},\boldsymbol{\Sigma}_{k}\right)$
$z_{i}^{(t)}$ $\sim\text{Categorical}\left(\pi_{1,t},\ldots,\pi_{K,t}\right),$
with observations independent across all particles $i$ and all time intervals $t$. Next, we model the mixture component means and probabilities as a regression on environmental covariates. However, instead of directly regressing on the principal components of the covariates, we regress on the hidden layer outputs of a random-weight neural network (Huang et al., 2006b, ):
$\boldsymbol{\mu}_{k,t}$ $=\boldsymbol{\beta}_{0,k}+{\boldsymbol{\beta}_{k}}^{\prime}\sigma\left(\boldsymbol{h}_{t}\right)$
$\pi_{k,t}$ $=\frac{\exp\left(\alpha_{0,k}+\sigma\left(\boldsymbol{h}_{t}\right)^{\prime}\boldsymbol{\alpha}_{k}\right)}{\sum_{\ell=1}^{K}\exp\left(\alpha_{0,\ell}+\sigma\left(\boldsymbol{h}_{t}\right)^{\prime}\boldsymbol{\alpha}_{\ell}\right)}.$
Here, $\boldsymbol{\beta}_{0,k}\in\mathbb{R}^{d\times 1}$, $\boldsymbol{\beta}_{k}\in\mathbb{R}^{n_{h}\times d}$, $\alpha_{0,k}\in\mathbb{R}^{1}$, and $\boldsymbol{\alpha}_{k}\in\mathbb{R}^{n_{h}}$ are regression coefficients. Also, $\sigma\left(\boldsymbol{h}_{t}\right)$ is a random-weight neural network with $n_{h}$ hidden nodes, constructed as follows. We first create a matrix $\boldsymbol{W}\in\mathbb{R}^{(p+1)\times n_{h}}$ of hidden layer weights and biases by randomly and independently drawing each element from a $\text{Unif}\left(-a,a\right)$ distribution, for a fixed constant $a$. Then, the principal components are transformed as
$\boldsymbol{h}_{t}=\boldsymbol{W}^{\prime}\begin{pmatrix}1\\ \boldsymbol{\Psi}_{t,\cdot}\end{pmatrix},$ (2)
where $\boldsymbol{\Psi}_{t,\cdot}$ is a column vector taken from row $t$ of $\boldsymbol{\Psi}$. Next, we apply a nonlinear activation function to extract nonlinear features. Specifically, denoting the elements of $\boldsymbol{h}_{t}$ as $h_{t,1},h_{t,2},\ldots,h_{t,n_{h}}$, we apply the softmax function (the element-wise logistic function):
$\sigma(\boldsymbol{h}_{t})=\left(\frac{1}{1+\exp(-h_{t,1})},\frac{1}{1+\exp(-h_{t,2})},\ldots,\frac{1}{1+\exp(-h_{t,n_{h}})}\right)^{\prime}.$ (3)
Crucially, the hidden weights are randomly drawn once and are fixed before the model training process, so the computational cost of estimating hidden layer weights via backpropagation is entirely avoided. Since a random-weight neural network can approximate any data-generating function with arbitrary accuracy (Huang et al., 2006a, ), it allows our model to learn nonlinear relationships between the response data’s mixture model parameters and the environmental covariates.
The approach taken in Hyun et al., (2023) is a special case of this model when $\boldsymbol{V}_{1:q,\cdot}=\boldsymbol{I}_{p}$, $\boldsymbol{W}=\boldsymbol{I}_{p+1}$, and the activation function $\sigma(\cdot)$ is set to be the identity function $\sigma(\boldsymbol{h}_{t})=\boldsymbol{h}_{t}$. In Sections 4 and 5 we compare our proposed model to one we call “linear”, which is identical to the one developed by Hyun et al., (2023) but uses principal components as regressors. This allows us to easily compare the interpretations of different models – nonlinear and linear – estimated using a common set of regressors. Also, using $q=9$ principal components enables simpler interpretation than if we had used the original $p=37$ covariates.
In all of our experiments, we set the hyperparamters $a=0.5$ and $n_{h}=70$. The $\mathrm{Unif}(-a,a)$ distribution used to draw the random weights in (2) only involves a single hyperparameter $a$, but can be replaced with any symmetric distribution centered at 0. We chose $a=0.5$ which induces roughly a $\mathrm{Unif}(0,1)$ distribution on $h_{t,1},\ldots,h_{t,70}$, and resulted in enough variation in the hidden layer weights to estimate a wide variety of regresssion functions. $n_{h}$ should be relatively large to allow flexibility in the model, although the marginal gain in accuracy diminshes at a certain point, and at the expense of increased computation. For choosing $n_{h}$ we recommend estimating models on a coarse $n_{h}$ grid (for example, $n_{h}=30,60,90,\ldots$), plotting the negative log-likelihood against $n_{h}$, and choosing $n_{h}$ as the value at which the negative log-likelihood starts to plateau.
3.1 Estimation
Following Hyun et al., (2023), we estimate model parameters by maximizing a constrained pseudolikelihood. Phytoplankton flow cytometry data naturally have a class imbalance due to the nature of small, lightweight cells to dominate in number. Larger organisms are less numerous but still account for a significant portion of the biomass in their habitat (Marañón,, 2015). Mixture models tend to overfit to the class which dominates in number (Xu and Jordan,, 1996), so we give additional importance to less numerous but more massive cells by exponentiating the likelihood contribution of each particle by its biomass. This results in the pseudolikelihood:
$\mathcal{L}(\alpha,\beta,\Sigma;\{\boldsymbol{y}_{i}^{(t)}\}_{i,t},\{\boldsymbol{x}_{t}\}_{t})=\prod_{t=1}^{T}\prod_{i=1}^{n_{t}}\left[\sum_{k=1}^{K}\pi_{k,t}(\boldsymbol{x}_{t})\phi_{d}\left(\boldsymbol{y}_{i}^{(t)}|\boldsymbol{\mu}_{k,t}(\boldsymbol{x}_{t}),\boldsymbol{\Sigma}_{k}\right)\right]^{c_{i}^{(t)}}.$ (4)
Here, we use $\alpha$ to denote the collection of coefficients $\{\alpha_{0,k}\}_{k},\{\boldsymbol{\alpha}_{k}\}_{k}$, and $\beta$ to denote $\{\boldsymbol{\beta}_{0,k}\}_{k},\{\boldsymbol{\beta}_{k}\}_{k}$. We also write $\pi_{k,t}(\boldsymbol{x}_{t})$ and $\boldsymbol{\mu}_{k,t}(\boldsymbol{x}_{t})$ to emphasize their dependence on the covariates. Pseudolikelihoods are commonly utilized in informative sampling scenarios, where the likelihood contributions are exponentiated by the observations’ inverse probabilities of selection. The resulting psuedolikelihood converges to the true population-generating distribution under certain regularity conditions (Savitsky and Toth,, 2016). In our dataset, particles with large biomass tend to be less numerous and therefore less likely to be observed, so it is reasonable to assume that a particle’s selection probability is approximately proportional to the inverse of its biomass.
In addition, Hyun et al., (2023) also impose two model constraints to encourage smoothness and penalize overly complex models, a technique we also utilize. First, since the optical properties of any particular species is limited in range due to biological constraints, we constrain the cluster means $\boldsymbol{\mu}_{k,t}$ to vary over time by up to a fixed amount $r$, called the maximum deviation. This constraint can be expressed as
$||\boldsymbol{\mu}_{k,t}-\boldsymbol{\beta}_{0,k}||_{2}=||\boldsymbol{\beta}_{k}^{\prime}\sigma(\boldsymbol{h}_{t})||_{2}\leq r,$
for cluster $k=1,\ldots,K$ and time $t=1,\ldots,T$. Another benefit of this constraint is that a properly chosen value of $r$ often prevents the switching of cluster labels over time. Second, to guard against selecting an overly complex regression model, lasso penalties (Tibshirani,, 2018) are imposed on all regression coefficients. After rescaling the pseudolikelihood, we arrive at the optimization problem
$\begin{split}\min_{\alpha,\beta,\Sigma}-\frac{1}{N}\mathcal{L}(\alpha,\beta,\Sigma;\{\boldsymbol{y}_{i}^{(t)}\}_{i,t},\{\boldsymbol{x_{t}}\}_{t})+\lambda_{\alpha}\sum_{k=1}^{K}||\boldsymbol{\alpha}_{k}||_{1}+\lambda_{\beta}\sum_{k=1}^{K}||\boldsymbol{\beta}_{k}||_{1}\\ \text{subject to }||\boldsymbol{\beta}_{k}^{\prime}\sigma(\boldsymbol{h}_{t})||_{2}\leq r\ \forall k=1,\ldots,K,\ \forall t=1,\ldots,T,\end{split}$ (5)
the solution of which provides parameter estimates. We solve (5) using the penalized EM algorithm described in Hyun et al., (2023) and use cross-validation to select the optimal hyperparameters $\lambda_{\alpha}$ and $\lambda_{\beta}$.
Lastly, since flow cytometry datasets have very large sample sizes $n_{t}$, we follow the approach taken by Hyun et al., (2023) to reduce computational demand by aggregating particles into a fixed grid of $D$ bins in each cytogram dimension. The new binned cytogram at time $t$ consists of observations $\left\{\boldsymbol{y}_{b}^{(t)}:b=1,\ldots,D^{d}\right\}$, where $\boldsymbol{y}_{b}^{(t)}$ is the center location of bin $b$. We associate with bin $b$ a new weight $c_{b}^{(t)}$ equal to the total biomass of particles within the bin, and use $\boldsymbol{y}_{b}^{(t)}$ and $c_{b}^{(t)}$ in place of $\boldsymbol{y}_{i}^{(t)}$ and $c_{i}^{(t)}$ in (5).
4 Simulation
4.1 Simulation design
In this section, we generate pseudo-synthetic data to test the numerical properties of our model. For each signal size $\Delta$ and at each time point $t=1,\cdots,296$, we generate $n_{t}=1,000$ particles $\boldsymbol{y}_{i}^{(t)}$ from a mixture of two one-dimensional Gaussian distributions. Specifically, the latent membership $z_{i}^{(t)}$ is first drawn from $\mathrm{Categorical}(\pi_{1,t},\pi_{2,t})$, and conditional on $z_{i}^{(t)}$, the particle’s value $\boldsymbol{y}_{i}^{(t)}\mid z_{i}^{(t)}=k$ is drawn from $\mathcal{N}\left(\mu_{1,t},0.2^{2}\right)$ if $k=1$, or $\mathcal{N}\left(\mu^{\Delta}_{2,t},0.2^{2}\right)$ if $k=2$. Cluster 1’s mean $\mu_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,4}$, taken from the real data’s principal component matrix $\boldsymbol{\Psi}$ in (1). In contrast, cluster 2’s mean $\mu_{2,t}^{\Delta}$ remains constant over time, so $\mu_{2,t}^{\Delta}=\mu_{2}^{\Delta}$ for all $t$. Also, cluster 1’s cluster probability $\pi_{1,t}$ changes over time $t$ as a function of the two principal components $\boldsymbol{\Psi}_{\cdot,1}$, and $\boldsymbol{\Psi}_{\cdot,2}$, and $\pi_{2,t}=1-\pi_{1,t}$ by construction.
There are three types of $\mu_{1,t}$ and $\pi_{1,t}$ configurations we consider:
  1. 1.
    Both linear. Cluster means and logit-transformed cluster probabilities are both linear functions.
  2. 2.
    Nonlinear probability. Fix the mean as a linear function and the logit-transformed cluster probabilities as nonlinear.
  3. 3.
    Nonlinear mean. Fix the logit-transformed cluster probabilities as linear and means as nonlinear.
For each configuration, we keep the first mean $\{\mu_{1,t}\}_{t=1}^{296}$ the same and vary the level of the second cluster mean $\mu_{2}^{\Delta}$ over signal sizes $\Delta$ which we define as the size of the gap between the two cluster means (averaged over time). That is, $\mu_{2}^{0}$ with the smallest signal size $\Delta=0$ has a zero gap between the average of the first cluster’s mean $\frac{1}{T}\sum_{t=1}^{T}\mu_{1,t}$ and $\mu_{2}^{\Delta}$. We consider equally spaced signal sizes between zero and the largest value $\Delta=0.95$, the signal size at which the two clusters’ ±$\pm 2$ standard deviation bands around the mean cease to overlap.
We use three types of nonlinearity: an interaction, a quadratic function, and a standard logistic function. The exact functions we use to generate the data are as follows (for $t=1,\cdots,296$):
  • Mean functions:
    • Linear: $\mu_{1,t}=1+0.015\boldsymbol{\Psi}_{t,1}+0.035\boldsymbol{\Psi}_{t,4}$
    • Nonlinear:
        Interaction: $\mu_{1,t}=1.05707+0.015\boldsymbol{\Psi}_{t,1}+0.02\boldsymbol{\Psi}_{t,4}+0.003\boldsymbol{\Psi}_{t,1}\boldsymbol{\Psi}_{t,4}$
        Quadratic: $\mu_{1,t}=1.030392+0.015\boldsymbol{\Psi}_{t,1}+0.01\boldsymbol{\Psi}_{t,4}|\boldsymbol{\Psi}_{t,4}|$
        Logistic:$\mu_{1,t}=1.089606-0.15(1+\exp(\boldsymbol{\Psi}_{t,1}))^{-1}+0.035\boldsymbol{\Psi}_{t,4}$
  • Probability functions:
    • Linear: $\text{logit}(\pi_{1,t})=-0.4\boldsymbol{\Psi}_{t,1}+0.1\boldsymbol{\Psi}_{t,2}$
    • Nonlinear:
        Interaction: $\text{logit}(\pi_{1,t})=-0.4\boldsymbol{\Psi}_{t,1}+0.02\boldsymbol{\Psi}_{t,1}\boldsymbol{\Psi}_{t,2}$
        Quadratic: $\text{logit}(\pi_{1,t})=-0.4\boldsymbol{\Psi}_{t,1}-0.05{\boldsymbol{\Psi}_{t,2}}^{2}$
        Logistic: $\text{logit}(\pi_{1,t})=-2+4.5(1+\exp(\boldsymbol{\Psi}_{t,1}))^{-1}+0.1\boldsymbol{\Psi}_{t,2}$.
These functions generate synthetic data which reproduce key characteristics of Prochlorococcus and PicoEukaryote observed in the Gradients 2 cruise, including realistic ranges of cell diameter and fluorescence intensities.
Next, for evaluating predictive out-of-sample accuracy, we use the log-pseudolikelihood in (4), calculated on an independently drawn dataset generated using the same parameters as the training dataset. For all scenarios, we use $5$-fold cross-validation to select tuning parameters, and for each model fit, we run the EM algorithm $30$ times. More specifically, we use a cross-validation scheme designed to prevent the data in the train and test folds from being too close to each other. Rather than using random splits of the time domain $\{1,\cdots,T\}$ for cross-validation folds, we partition it into blocks of 20 consecutive time intervals and then distribute the blocks among the folds. Specifically, for $K=5$ folds, the first fold is assigned blocks $1,6,11,\cdots$ and contain time points $\{1,\cdots,20\}\cup\{121,\cdots,140\}\cup\cdots$, the second fold is assigned the blocks $2,7,12,\cdots$, and so forth.
Figure 2: Estimated model on simulated 1-dimensional data. Both panels display in the background the simulated and binned data $y_{b}^{(t)}$ over time $t$, generated from the “Interaction Mean” model, with signal size $\Delta=0.40384$. The greyscale color of the bin $b$ at time $t$ is proportional to the biomass $c_{b}^{(t)}$. Overlaid are the estimated parameters $\hat{\boldsymbol{\mu}}_{k,t}$ over time $t$ from the linear (left) and nonlinear (right) models as solid colored lines. These lines’ thickness at time $t$ is proportional to the estimated cluster probability $\hat{\pi}_{k,t}$. The thin colored lines are the bounds of symmetric 95% pointwise conditional probability regions of each Gaussian. The black dashed lines represent the true cluster means, and the black dotted lines represent the true 95% central probability regions.
4.2 Simulation results
Figure 2 shows an example of the estimated linear model (left panel) and nonlinear model (right panel), both using data generated from a linear probability function and an interaction mean function. In the true model, the variation in the first cluster mean $\{\mu_{1,t}\}_{t=1}^{296}$ over time is largely due to $\boldsymbol{\Psi}_{\cdot,4}$, which is highly correlated with sunlight and thus oscillates regularly with a daily periodicity. The interaction in the linear effect of $\boldsymbol{\Psi}_{\cdot,4}$ with $\boldsymbol{\Psi}_{\cdot,1}$ causes the true amplitude of the oscillation to become smaller at high latitudes. However, the estimated linear model fails to capture this and the amplitude of the estimated oscillation in ${\mu}_{1,t}$ does not shrink enough in high latitudes. It also noticeably underestimates the first cluster mean $\mu_{1,t}$ when $t\in[40,110]$, as well as the second cluster mean $\mu_{2,t}$ across all time points. In contrast, the nonlinear model more accurately captures the change in amplitude, with its estimated means closely matching the ground truth at all time points. Figure 3 shows the aggregate summary of the simulation results. The y-axis values in each panel is the negative log-pseudolikelihood (NLPL) of an estimated model after subtracting the NLPL model of the oracle, or true generating model. It is clear that the nonlinear model almost uniformly outperforms the linear model in terms of the out-of-sample negative log-pseudolikelihood, which demonstrates a better fit to the data. (The nonlinear model even outperforms for linear data sometimes, which may be due to large sample sizes with relatively little noise, which makes training and test datasets very similar.)
Figure 3: Out-of-sample negative log-pseudolikelihoods (NLPLs) after subtracting the NLPL of the oracle of each model at each signal size. The linear model is represented by the dashed line and the nonlinear by the solid line.
5 Application to SeaFlow cruise data
In this section, we estimate both the linear model and the nonlinear model on the real flow cytometry dataset described in Section 2. We bin the cytograms $\boldsymbol{y}^{(t)}$ with a grid size of $D=40$ bins in each dimension and set bin weights $c_{b}^{(t)}$ equal to each bin’s total biomass. We followed the same cross-validation scheme as described in Section 4 with $10$ EM algorithm restarts.
Figure 4: Cytogram in two dimensions, diameter and chlorophyll, at time $t=33$ (2017-07-02 08:00-09:00), with model estimates overlaid in red (marking the important clusters) or grey (the rest). The data in the background is binned, with the intensity of blue hue proportional to the total biomass in each bin. The solid points mark the cluster means, and the size of the points are proportional to the estimated cluster probabilities. The ellipses represent 2-dimensional projections of symmetric cluster-conditional 95% probability sets of each cluster.
The models’ estimates at the time point $t=33$ are shown in Figure 4, where we focus on two out of the three cytogram dimensions, and overlay the estimated model with the binned data. The resulting clusters from each model have subsequently been expert-identified as Prochlorococcus (Pro), Synechococcus (Syn), and two distinct PicoEukaryote (Pico1 and Pico2), as well as other populations. It is visually clear that the two models identify the major populations (Pro, Syn, Pico1 and Pico2 in Figure 4) similarly.
We also measured the out-of-sample accuracy of each model, using the following nested cross-validation scheme:
  1. 1.
    Partition the time points $\mathcal{T}=\{1,\ldots,T\}$ into folds $\mathcal{T}^{(j)}$, $j=1,\ldots,5$, using equally spaced blocks of 20 consecutive time points. Partition the data accordingly as $\mathcal{D}^{(j)}=\left\{\left(\boldsymbol{y}_{i}^{(t)},\boldsymbol{x}_{t}\right):t\in\mathcal{T}^{(j)}\right\}$, for fold $j=1,\ldots,5$.
  2. 2.
    For outer fold $j=1,\ldots,5$:
    1. (a)
      Use $\mathcal{D}^{(j)}$ as a held-out dataset. Using the remaining data $\bigcup_{k\neq j}\mathcal{D}^{(k)}$, cross-validate the model to select the optimal hyperparameters, then refit the model as described in Section 3.1, resulting in model parameter estimates $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\Sigma}$.
    2. (b)
      Evaluate the unpenalized negative log-pseudolikelihood $-\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$.
  3. 3.
    Let $-\frac{1}{5}\sum_{j=1}^{5}\mathcal{L}\left(\hat{\alpha},\hat{\beta},\hat{\Sigma};\{\boldsymbol{y}_{i}^{(t)}\}_{i,t\in\mathcal{T}^{(j)}},\{\boldsymbol{x}_{t}\}_{t\in\mathcal{T}^{(j)}}\right)$ be the final estimate of the out-of-sample model fit.
The linear model had an out-of-sample negative log-pseudolikelihood of $3.446$ while the nonlinear model’s was $3.439$, showing similar generalization performance. In addition, the relative abundance predictions of the two models are very similar, as Figure 5 shows. However, there are notable differences between the two models’ cluster mean predictions. The linear model predicts a dramatic decrease in Prochlorococcus diameter between July 2nd and July 6th, while the nonlinear model predictions remain relatively stable, showing a diel oscillation at a similar overall level. Also, the nonlinear model better captures the amplitude of the expected diel oscillation in chlorophyll of Prochlorococcus throughout all time points, matching its expected behavior from the corresponding oscillation in cell size. Overall, while the models display similar clustering results, the nonlinear model estimates more flexible cluster mean movement over time that fits scientific intuition.
Figure 5: Mean and probability estimates over time for the Prochlorococcus cluster. In the left panel, the three dimensions of the cluster mean are plotted together; diameter, chlorophyll, and phycoerythrin estimates are represented by the red, blue, and yellow lines, respectively. In the right panel, relative abundance predictions are displayed. The nonlinear model predictions are represented by the solid lines and the linear model predictions by dot-dashed lines. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone.
Next, we interpret the estimated models in relation to the principal components of the environmental covariates. Since we use a neural network with random weights, the parametric form of the estimated regression equations is difficult to interpret analytically. Instead, we interpret the relationship between the parameters $(\boldsymbol{\mu}_{k,t},\pi_{k,t})$ and principal components $\boldsymbol{\Psi}_{t,j}$ visually using partial response curves. We define partial response curves as the average of a model parameter estimate given a particular value of principal component $j$ and average values of all other principal components, constructed as follows. For example, let us denote as:
$\hat{\boldsymbol{\mu}}_{k}(\psi_{1},\cdots,\psi_{q})=\boldsymbol{\hat{\beta}}_{0,k}+\boldsymbol{\hat{\beta}}_{k}^{\prime}\sigma(\boldsymbol{W}^{\prime}\left(\begin{smallmatrix}1\\ \psi_{1}\\ \vdots\\ \psi_{q}\end{smallmatrix}\right)),$
the $k$th cluster mean predictions evaluated at scalar principal component values $\psi_{1},\cdots,\psi_{q}$. Also denote $\bar{\boldsymbol{\Psi}}_{\cdot,l}$ as the time-wise average of the $l$th principal component. Then, the partial response curve of the $k$th cluster mean $\boldsymbol{\mu}_{k}$ in response to changes in the $j$th principal component is represented as 2-dimensional points:
$(x,\;\hat{\boldsymbol{\mu}}_{k}(\bar{\boldsymbol{\Psi}}_{\cdot,1},\cdots,\bar{\boldsymbol{\Psi}}_{\cdot,j-1},x,\bar{\boldsymbol{\Psi}}_{\cdot,j+1},\cdots,\bar{\boldsymbol{\Psi}}_{\cdot,q})),$
calculated over a fine grid of values of $x$ in the range of $j$th principal component values $\boldsymbol{\Psi}_{\cdot,j}$ in the data. Partial response curves for relative abundance $\pi_{k}$ are defined similarly.
Figure 6: Partial responses of Prochlorococcus diameter, the first PicoEukaryote cluster’s log diameter, and the second PicoEukaryote cluster’s relative abundance as functions of the first principal component. The vertical dotted lines represent the boundaries of the North Pacific Transition Zone.
Next, as a visual aid for interpreting the results, we plot two vertical dotted lines in each panel of Figure 5 and Figure 6 representing the approximate start and end of the North Pacific Transition Zone (NPTZ), estimated directly from the data (bottom-left panel of Figure 1). These lines mark a region where ocean conditions rapidly change latitudinally.
We see in Figure 6 a few representative partial response plots in which certain phytoplankton properties are displayed as functions of PC1. As PC1 increases, the cruise moves northward closer to the subarctic region. We see in the left panel of Figure 6 that both models predict Prochlorococcus diameter to be larger monotonically as one moves to higher latitudes, if one were to hold as constant ocean water conditions that are not associated with latitude (PC1). This strongly suggests that the average Prochlorococcus microbe is larger in higher latitudes. However, the linear model predicts only a flatter and linear change, by design. The nonlinear model is capable of predicting a more flexible change. It predicts Prochlorococcus diameter to change (1) much slower in the south-most and north-most regions along the cruise track, and (2) much faster in the transition zone, especially within the NPTZ. This is consistent with external studies (Thompson et al.,, 2021) that different subtypes of Prochlorococcus are dominant in each of the two gyres – Subpolar and Subtropical – and our model estimates further suggest that those subtypes have a clear size difference.
For PicoEukaryote, both models predict that the cell diameter decreases monotonically as one goes from south to north along the cruise trajectory. As with Prochlorococcus, the cell diameter levels off to become flat for high latitudes, indicating stable populations of PicoEukaryote at high latitudes. In fact, when PC1 approaches ±$\pm 8$, both PicoEukaryote and Prochlorococcus clusters’ cell diameters are predicted to flatten, reflecting the existence of maximum and minimum sizes of plankton cells. The linear model predictions, however, do not level off, and any prediction outside of the observed latitude range is sure to be too extreme to be scientifically useful.
Finally, the second PicoEukaryote cluster (Pico2) is predicted to have a maximum relative abundance at a value of PC1 which corresponds to latitude 38.8 N. At more northern locations than this, the relative abundance is predicted to decrease. Such results highlight the possibility that nonlinear mixture-of-experts models, such as our proposed model, can be used to learn globally and locally optimal ocean conditions where certain phytoplankton species are most relatively abundant.
Figure 7: The top panel displays partial response curves of certain model parameters as functions of the second principal component, conditional upon several different values of the first principal component. The bottom panel displays the trajectory of the Gradients 2 cruise, with horizontal lines marking the values of PC1 used in the top panel as well as the NPTZ boundaries.
The nonlinear model also captures interaction effects between PC1 and PC2, as displayed in Figure 7. In each panel, the three lines show a conditional partial response curve of the nonlinear model’s parameter, say $\theta$, defined by first fixing PC1 to be a certain value $c$, then plotting the partial response of the parameter estimate $\hat{\theta}$:
$(x,\;\hat{\theta}(c,x,\bar{\boldsymbol{\Psi}}_{\cdot,3},\cdots,\bar{\boldsymbol{\Psi}}_{\cdot,q}))$
over a fine grid of values of $x$ in the range of the second principal component $\boldsymbol{\Psi}_{\cdot,2}$. Figure 7 highlights the conditional partial responses of the cluster probability $\hat{\pi}_{k,t}$ (in the left panel) or one dimension of the cluster mean $\hat{\boldsymbol{\mu}}_{k,t}$ (in the center and right panel), fixing the values of PC1 at three values ($-6$, $-1$, and $4$) which correspond to three locations clearly south of, inside of, and north of the North Pacific Transition Zone.
First, for cluster means, whenever the three lines from the same model are not parallel and equidistant, this indicates an interaction between PC1 and PC2. We see this most clearly in PicoEukaryote diameter in the right panel. At a high value of PC1 (4, corresponding to north of the NPTZ), the partial response curve is nearly linear. However, at lower values of PC1 (inside or south of the NPTZ), the partial response to PC2 clearly becomes nonlinear. In the center panel showing the phycoerythrin fluorescence of Synechococcus, the partial response curve changes from convex for lower PC1 values to concave for higher PC1 values.
Next, we examine the relative abundance prediction of Prochlorococcus (left panel of Figure 7). First, the linear model’s partial response curve to PC2 follows a similar logistic shape at all specified values of PC1. On the other hand, the nonlinear model curve clearly changes in shape, especially from PC1 = $-1$ to PC1 = $-6$. This suggests an interaction effect between PC1 and PC2 which allows the response of Prochlorococcus relative abundance to PC2 to be much more flexible depending on the latitude, of which PC1 serves as a proxy. Furthermore, the nonlinear model has a shallower slope overall, thus estimating Prochlorococcus abundance to be more stable in response to physical oceanographic attributes (PC2) in southern waters than the linear model does.
6 Discussion
In this article we propose a novel mixture of neural network experts which uses randomly generated and fixed hidden layer weights to represent a nonlinear response to the covariates. This allows one to capture nonlinear behavior in flow cytometry datasets with the key advantage of a relatively low computational cost. We also use principal components as regressors to allow for compression of environmental covariates. When paired with partial response plots, using principal component regressors enhances our ability to interpret the model with respect to major modes of variation in the environmental factors. We conduct an extensive simulation to show that the proposed model matches the clustering performance of similar linear models and more accurately fits nonlinear data. Also, we apply our model to real phytoplankton flow cytometry data collected along a north-south transect on the North Pacific Ocean. The results show that the nonlinear model better reflects expected behaviors of phytoplankton traits when compared to the linear model. For example, the nonlinear model predicts the rate of change in cells’ optical properties and abundances to depend on how rapidly the environmental conditions change. Furthermore, optical properties are predicted to level off when environmental conditions are at very low or high values, consistent with the fact that microbes of a certain species have a biological upper and lower limit to their size. This saturation behavior represents a fundamental biological constraint that linear models cannot capture, as they assume constant rates of change across the entire environmental gradient. The nonlinear approach allows different expert components to become active at different portions of the gradient, enabling the model to capture rapid responses under moderate conditions while transitioning to asymptotic behavior at environmental extremes. We therefore recommend the use of a nonlinear model when analyzing flow cytometry samples collected over a range of locations with diverse environmental conditions.
There is room for further model development. Our method assumes clusters are Gaussian, yet cell clusters in cytograms often exhibit skewness and heavy tails. A Bayesian semiparametric extension may provide a better fit to the data with the added benefits of uncertainty quantification and hyperparameter tuning without resampling procedures or cross-validation. Another area for future work relates to the hardware limitations of flow cytometers; the Gradients 2 dataset contains some observations which are censored at SeaFlow’s measurement limits. Proper handling of censored data may strengthen model validity and accuracy. Finally, it would be valuable to extend our application to a larger collection of cruise datasets exploring an even wider range of locations and times. Mapping the spatial distributions and average optical properties of phytoplankton regionally or globally based on in situ data, uncovering temporal trends, and identifying the environmental conditions which are most predictive of phytoplankton behavior, are ambitious yet exciting possibilities.
Acknowledgements
This work was supported through grants by the Simons Collaboration on Computational Biogeochemical Modeling of Marine Ecosystems/CBIOMES (Grant ID: 1195553 to Sangwon Hyun) and the Simons Microbial Oceanography Project Award (Grant ID: 574495 to Francois Ribalet) and the National Science Foundation award NCSE-2215169. The authors also acknowledge the Hummingbird Computational Cluster at the University of California, Santa Cruz, for providing computing resources that have contributed to the research results reported within this publication. URL: https://hummingbird.ucsc.edu/.
Supplement A: Principal Components over Time
Figure 8: Principal component 2 and its most important constituent covariates over time: iron (Fe), sea level anomaly (SLA) and wind speed.
Figure 9: All nine principal components used in our data analysis.
Supplement B: SeaFlow Cruise Data Clustering Results
Figure 10: Estimated means and probabilities over time of Synechococcus for each model.
Figure 11: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 1) for each model.
Figure 12: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 2) for each model.
Supplement C: Individual Conditional Expectations
Figure 13: Partial response curves of each response variable for each species (Pro, Syn, Pico1, and Pico2) as functions of the second principal component.
Figure 14: Partial response curves of each response variable for each species as functions of the second principal component.